OpenStudy (anonymous):

lne^3 evaluating logarithms

OpenStudy (anonymous):

ln(E^x)= x

OpenStudy (anonymous):

by definition

OpenStudy (amorfide):

use your laws of natural logarithms \[lna^{n}=n \times lna\] when a=e and n=3 \[lne^{3}=3lne\] since natural log is base e base e of e, is the same as saying \[e^{n}=e\] therefore lne=1 this gives \[3lne=3 \times 1\]